# Michael William Healy-Kalesh Association between Recurrent Novae and Nova Super-Remnants

### v2

@@ 3 Predicting a Population of Nova Super-Remnants
Regardless of the combination of parameters used, all modelled RNe grew a NSR (as shown in Figure 2), including novae with long inter-eruption periods and even those systems with an eroding WD (opening up the possibility of NSRs existing around old novae with low mass WDs). To reiterate this key finding: all novae should be surrounded by a NSR [66]. Though, only NSRs hosted by novae with high accretion rates will be currently observable. Akin to Darnley et al. [62], Healy-Kalesh et al. [66] found that nova systems with high mass accretion rates () can create NSRs (up to parsecs across) consisting of a very low density cavity surrounded by a hot ejecta pile-up region, all contained within a thin, high density shell [66]. Though, less energetic ejecta combined with longer intervals early in the evolution of the Healy-Kalesh et al. [66] runs (allowing for prolonged phases of radiative cooling and energy loss) significantly altered the structure of the remnant shells compared to the NSR shell in Darnley et al. [62], such that they were substantially thinner and denser. For the range of parameters considered, higher ISM densities and higher both restricted the radial size of the growing NSR whereas the temperature and initial mass of the WD both had a relatively small influence on the evolution of the NSR [66]. Notably, NSRs grown from oxygen-neon (ONe) WDs (initial WD mass ) were found to be considerably reduced in size (see bottom row of Figure 2), making these types of NSRs more difficult to find.
@@ 5 RS Ophiuchi Nova Super-Remnant
With RS Ophiuchi being a Galactic RN, and arguably the most well-studied nova after 12a, it represents another worthy candidate for hosting a NSR. Modelling from Healy-Kalesh et al. [66] demonstrated that, for an estimated ISM density, the potential NSR shell around RS Oph would have dimensions of pc with a cavity pc, equating to and , respectively [70]. As the cavity exists due to an almost complete removal of all cool gas from the inner regions of the NSR, a search through near-infrared IRIS (Improved Reprocessing of the IRAS Survey) data [71] proved to be fruitful. A cavity-like feature was promptly found with dimensions (if co-located with RS Oph) of pc, matching the predicted cavity size [66], and so was claimed to be likely associated with the central RN [70]; the same feature was independently identified over a decade before [72], though the wind from RS Oph was suggested to be responsible for its creation.
Instead, the claim of a NSR cavity [70] was given further credence through the independent discovery of a H-bright NSR shell. With the success of unveiling the second NSR around KT Eri, Condor was used to target RS Oph. Observing the surroundings of RS Oph with lengthy exposures revealed another vast structure, again bright in H and [N II] emission. At a distance of RS Oph, the uncovered nebulosity measures 30 pc across and so takes its place as the third nova super-remnant discovered (Michael Shara, presented during this conference and private communication). Not only did this discovery of a H-bright NSR shell around RS Oph reinforce the connection between RNe and NSRs but it also complimented and ultimately strengthened the separate discovery of the NSR-like feature found within archival near-IR imaging presented in Healy-Kalesh et al. [70].

### v4

@@ 3 Predicting a Population of Nova Super-Remnants
Regardless of the combination of parameters used, all modelled RNe grew a NSR (as shown in Figure 2), including novae with long inter-eruption periods and even those systems with an eroding WD (opening up the possibility of NSRs existing around old novae with low mass WDs). To reiterate this key finding: all novae should be surrounded by a NSR [66]. Though, only NSRs hosted by novae with high accretion rates will be currently observable. Akin to Darnley et al. [62], Healy-Kalesh et al. [66] found that nova systems with high mass accretion rates () can create NSRs (up to parsecs in radius) consisting of a very low density cavity surrounded by a hot ejecta pile-up region, all contained within a thin, high density shell [66]. Though, less energetic ejecta combined with longer intervals early in the evolution of the Healy-Kalesh et al. [66] runs (allowing for prolonged phases of radiative cooling and energy loss) significantly altered the structure of the remnant shells compared to the NSR shell in Darnley et al. [62], such that they were substantially thinner and denser. For the range of parameters considered, higher ISM densities and higher both restricted the radial size of the growing NSR whereas the temperature and initial mass of the WD both had a relatively small influence on the evolution of the NSR [66]. Notably, NSRs grown from oxygen-neon (ONe) WDs (initial WD mass ) were found to be considerably reduced in size (see bottom row of Figure 2), making these types of NSRs more difficult to find.
@@ 5 RS Ophiuchi Nova Super-Remnant
With RS Ophiuchi being a Galactic RN, and arguably the most well-studied nova after 12a, it represents another worthy candidate for hosting a NSR. Modelling from Healy-Kalesh et al. [66] demonstrated that, for an estimated ISM density, the potential NSR shell around RS Oph would have dimensions of pc with a cavity pc, equating to and , respectively [70]. As the cavity exists due to an almost complete removal of all cool gas from the inner regions of the NSR, a search through far-infrared IRIS (Improved Reprocessing of the IRAS Survey) data [71] proved to be fruitful. A cavity-like feature was promptly found with dimensions (if co-located with RS Oph) of pc, matching the predicted cavity size [66], and so was claimed to be likely associated with the central RN [70]; the same feature was independently identified over a decade before [72], though the wind from RS Oph was suggested to be responsible for its creation.
Instead, the claim of a NSR cavity [70] was given further credence through the independent discovery of a H-bright NSR shell. With the success of unveiling the second NSR around KT Eri, Condor was used to target RS Oph. Observing the surroundings of RS Oph with lengthy exposures revealed another vast structure, again bright in H and [N II] emission. At a distance of RS Oph, the uncovered nebulosity measures 30 pc across and so takes its place as the third nova super-remnant discovered (Michael Shara, presented during this conference and private communication). Not only did this discovery of a H-bright NSR shell around RS Oph reinforce the connection between RNe and NSRs but it also complimented and ultimately strengthened the separate discovery of the NSR-like feature found within archival far-IR imaging presented in Healy-Kalesh et al. [70].